OpenStudy (anonymous):

The cook needs to make the dough for 9 small pizzas. Each pizza dough needs 1 1/3 cup of flour. How much flour does the cook need to make all the pizza dough?

OpenStudy (anonymous):

Just multiply them. I'll set it up and you multiply. \[\frac{ 9 }{ 1 }*\frac{ 4 }{ 3 }=?\]

OpenStudy (anonymous):

3/36?

OpenStudy (anonymous):

The opposite. It would be \[\frac{ 36 }{ 3 }=12\]